OpenStudy (samigupta8):

Given that f'(x)= product of (n-x)^n from n=1 to 101. No. Of relative maximum of f(x) is

OpenStudy (samigupta8):

@kainui

OpenStudy (samigupta8):

Product is like (1-x)(2-x)^2(3-x)^3 ....and so on

OpenStudy (samigupta8):

@hartnn

OpenStudy (samigupta8):

@parthkohli

Parth (parthkohli):

Only studied this in physics. 1. \(f'(x_0)=0\) 2. \(f''(x_0) < 0\)

OpenStudy (samigupta8):

Yep it is this way only in maths too... This is second derivative test though..

Parth (parthkohli):

Yeah, what else are we supposed to do?

OpenStudy (samigupta8):

First derivative says that if f'(x) changes sign from positive to negative then there occurs a maxima And reverse for minima

OpenStudy (samigupta8):

@michele_laino

OpenStudy (michele_laino):

I think that the sign of the first derivative is the sign of this function: \[g\left( x \right) = \left( {1 - x} \right)\left( {2 - x} \right)\left( {3 - x} \right)...\left( {101 - x} \right)\]

OpenStudy (samigupta8):

Sir how ? Actually we can do one thing Since (2-x)^2 is all time positive similarly such terms like (4-x)^4 is also +ve

OpenStudy (samigupta8):

So we can ignore them .. Isn't it?

OpenStudy (samigupta8):

Then we will be left with only odd term which again will include just (1-x)(3-x)(5-x)...and so on.

OpenStudy (michele_laino):

sorry I have made a typo, there are odd numbers only: \[g\left( x \right) = \left( {1 - x} \right)\left( {3 - x} \right)\left( {5 - x} \right)\left( {7 - x} \right)...\left( {101 - x} \right)\]

OpenStudy (samigupta8):

Yes ...sir it has to be like this now...

OpenStudy (michele_laino):

\(g(x)\) is positive when \(all\) factors are positive, since we have an \(odd\) number of factors

OpenStudy (michele_laino):

or when there are an \(even\) number of negative factors

OpenStudy (samigupta8):

BT sir there are even no. of factors

OpenStudy (samigupta8):

From 1 to 101 we have even no. of terms in this case....

OpenStudy (michele_laino):

yes! you are right! sorry again!

OpenStudy (samigupta8):

No problem...

OpenStudy (samigupta8):

Sir what next?

OpenStudy (samigupta8):

If we have even no of terms then we can say that product will be positive when they are pairwise positive or negative

OpenStudy (michele_laino):

for example if we consider these three factors: (I'm applying the principle of Mathematical induction) \[g\left( x \right) = \left( {1 - x} \right)\left( {3 - x} \right)\left( {5 - x} \right)\]

OpenStudy (michele_laino):

we have this diagram:  where the sign refer to the sign of product of \((1-x)(3-x)(5-x)\)

OpenStudy (michele_laino):

as we can see only he points \(x=1,\;x=5\) are maximum for \(g(x)\)

OpenStudy (michele_laino):

the*

OpenStudy (michele_laino):

so, by means of the principle of Mathematical induction, wehave the maxi8mum points at these ponts: \(x=1,5,9,13,...\) namely the points of maximum for \(f(x)\) are aan arithmetic sequence, whose first term is \(1\), and the constant is \(4\), and the last term is \(x=101\) so we have: \[101 = 1 + \left( {n - 1} \right) \times 4\] from such equation, I get \(n=26\) points of maximum

OpenStudy (samigupta8):

Thank you so much sir!!!